How many different samples of size n=3 can be obtained from a finite population of size N=9 . 

Select one: a. 42 b. 84 c. 21 d.7 


1
Expert's answer
2021-11-01T11:47:13-0400

Given: "n=3, N=9"

Use the formula for combinations

"C (n, r) = \frac{n!}{(n-r!).r!}"

Substitute the given in each letter, where in this question, given population N is for n in the formula, and the given sample size n is for r in the formula.

"C(9,3)=\frac{9!}{(9-3!).3!}
\\=\frac{9!}{6!.3!}
\\=\frac{9.8.7.6!}{6!.3.2}
\\=84"

Option (b) is correct.